OpenStudy (anonymous):

the resultant of two forces of magnitude 10N each has also magnitude of 10N.the angle between the forces is? (a) o (b)90 (c)180 (d)120

OpenStudy (anonymous):

It is an equal lateral triangle therefore each angle is equal to 60 degrees. So does that mean the total of the three which means 180?

OpenStudy (anonymous):

the angle between them will be 120.Let me explain .resultant of two vectors A and B is given as: R=\[\sqrt{A ^{2}+B ^{2}+2ABcos \theta}\] cos 120=-1/2.so in this case, R=\[\sqrt{10^{2}+10^{2}-2\times\ 10 \times10\times1/2}\].which gives R=10.